(a) What is the maximum angular magnification of an eyeglass lens having a focal length of 19.6 cm when used as a simple magnifier?


(b) What is the magnification of this lens when the eye is relaxed?
Answer #1

Part 1)
The maximum angular magni?cation of a lens having a focal length of 19.6 cm is

(25cm/ f) + 1= 25 cm/19.6cm +1 = 1.27+ 1 =2.27

f = 11.01 cm

Part 2)
The angular magni?cation of this lens when the eye is relaxed is
(25cm/ f) = 1.27